### v2

@@ 5 Budget-adaptivity Gap
Corollary 4.
The approximation factor of the greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}$ is at least $1-\frac{1}{e}$, for any $\delta,\xi>0$.
@@
Assume by contradiction that the approximation factor $\alpha$ of ${\sf MultiGr}_{\delta,\xi}$ is higher than $1-\frac{1}{e}$, for some $\delta,\xi>0$. Let $\epsilon:=\frac{e}{e-1}-\frac{1}{\alpha}>0$ be a fixed constant, $I(\epsilon)$ be the input instance considered in the proof of Theorem 2, $\pi^{*}$ and $\pi$ be respectively an optimal fully adaptive policy and an optimal partially adaptive policy for $I(\epsilon)$. In Theorem 2 we showed that the budget adaptivity gap of $I(\epsilon)$ is strictly higher than $\frac{e}{e-1}-\epsilon=\frac{1}{\alpha}$. Then, as ${\sf MultiGr}_{\delta,\xi}$ is a partially adaptive greedy policy, we necessarily have that $\sigma(\pi)\geq\sigma({\sf MultiGr}_{\delta,\xi}(I(\epsilon)))$. We conclude that $\alpha>\frac{1}{BGap(I(\epsilon))}=\frac{\sigma(\pi)}{\sigma(\pi^{*})}\geq\frac{\sigma({\sf MultiGr}_{\delta,\xi}(I(\epsilon)))}{\sigma(\pi^{*})},$ but this contradicts the $\alpha$-approximation assumed at the beginning.
6 Conclusions
In this paper we presented the first results for the problem of adaptively optimizing the value of a stochastic submodular function over multiple rounds with total budget constraints when the state distributions are known.
Our analysis aims to provide theoretical guarantees on the approximation that can be achieved in polynomial time. We proposed two algorithms: one based on a dynamic programming approach, that is able to compute the optimal multi-round adaptive policy in polynomial time whenever a single-round optimization problem associated with each round can be efficiently solved (e.g., if the number of items is constant and the number of random realizations is polynomial), and a simpler greedy algorithm that, despite being only partially adaptive, still returns a $(1-1/e-\epsilon)$-approximation of the optimum, and such result is essentially optimal in light of the hardness result of [23] (holding for simpler non-adaptive single-round maximization problems). Furthermore, by introducing the budget-adaptivity gap, we measured how much is lost, in terms of value obtained, if we distribute the budget over all rounds in advance, i.e. in a non-adaptive way.
A first research direction is that of combining multiple rounds, budget and adaptivity in other classical optimization problems, as we did in submodular maximization problems, and then study computational complexity, approximation and budget-adaptivity gap.
Another interesting open question is related to the efficiency of the above greedy algorithm in terms of run-time. Indeed, even if the proposed approximation algorithm we provided is essentially simple and it is proved to run in polynomial time, it involve some routines, such as Monte Carlo simulations, that are known to limit their practical application. Anyway, we believe that, as occurred for many similar problems, our algorithm may foster further research towards more practical approaches, that, e.g., replace the more expensive routines with heuristics. Not only, but our algorithm will also serve as a benchmark for the evaluation of these heuristics.

### v3

@@ D.2 Full Proof of Theorem 2 (for approximated oracles)
$\begin{split}&\sum_{t\in[T]}\left(1-\frac{1}{e}\right)OPT_{t}(b_{t}^{*})-(\delta+\xi\Lambda)B=\sum_{t\in[T]}\left(\left(1-\frac{1}{e}\right)OPT_{t}(b_{t}^{*})-(\delta+\xi\Lambda)b_{t}^{*}\right)\\ &\quad\leq\sum_{t\in[T]}{GR}_{t}(b^{*}_{t})={GR}(\vec{b}^{*}).\end{split}$ (48)
By exploiting the greedy assignment of ${\sf BudgetGr}_{\delta,\xi}$ and the adaptive submodularity, we have
$GR(\vec{b}^{*})-(3\delta+2\xi\Lambda)B\leq GR(\vec{b})$ (49)
(see Lemma 6 below). Furthermore, we have
$(4\delta+3\xi\Lambda)B\leq\epsilon\cdot\lambda\leq\epsilon\cdot OPT,$ (50)
where the first inequality holds by definition of $\delta$ and $\xi$, and the last one holds by Lemma 6 below. By combining inequalities (4-7) and (48-50) we obtain
$\frac{1}{2}\left(1-\frac{1}{e}-\epsilon\right)OPT\leq\left(1-\frac{1}{e}\right)\sum_{t\in[T]}OPT_{t}(b^{*}_{t})-\epsilon\cdot OPT\leq\left(1-\frac{1}{e}\right)\sum_{t\in[T]}OPT_{t}(b^{*}_{t})-(4\delta+3\xi\Lambda)B$
$\quad\leq GR(\vec{b}^{*})-(3\delta+2\xi\Lambda)B\leq GR(\vec{b}),$
and this shows the claim.
D.3 Technical Lemmas for the Full Proof of Theorem 2 under Approximated Oracles
Lemma 5.
For any $t\in[T]$ and integer $b\geq 0$, we have
$\left(1-\frac{1}{e}\right)OPT_{t}(b)-(\delta+\xi\Lambda)b\leq GR_{t}(b).$
@@ D.3 Technical Lemmas for the Full Proof of Theorem 2 under Approximated Oracles
Fix $t\in[T]$ and an integer $b\geq 0$. For any $i\in[b]$, let $GR^{*}_{t}(i-1,b)$ be the value of an adaptive policy $\pi^{*}_{t}(i-1,b)$ at round $t$ that first executes the greedy single-round policy ${\sf SingleGr}_{\delta,\xi}$ at round $t$ with budget $i-1$, then runs an optimal adaptive policy $\pi^{*}_{t}$ at round $t$ with budget $b$, without observing the realizations coming from the greedy one, and finally selects the items previously chosen by both policies. By the monotonicity property, we have
$OPT_{t}(b)\leq GR^{*}_{t}(i-1,b).$ (51)
Furthermore, by exploiting the adaptive submodularity, we can show that inequality
$GR_{t}(i)\geq\frac{GR^{*}_{t}(i-1,b)}{b}+\left(1-\frac{1}{b}\right)GR_{t}(i-1)-(\delta+\xi\Lambda)$ (52)
holds for any $i\in[b]$. Indeed, let $\bm{v}_{t,i}$ denote the (random) item selected by the greedy single-round policy ${\sf SingleGr}_{\delta,\xi}$ at round $t$ and step $i$, $\bm{S}_{t,i-1}:=\{\bm{v}_{t,1},\ldots,\bm{v}_{t,i-1}\}$ denote the set of the first $i-1$ items selected by ${\sf SingleGr}_{\delta,\xi}$, and $\Delta_{t,i}:=GR_{t}(i)-GR_{t}(i-1)$ denote the expected increment of $f_{t}$ achieved at step $i$ of the greedy single-round policy. Furthermore, for any partial state $\langle S_{t,i-1}\rangle\sim\langle\bm{S}_{t,i-1}\rangle$, let $A(\langle S_{t,i-1}\rangle)$ denote the event “$\max_{v\in V}\Delta_{t}(v|\langle S_{t,i-1}\rangle)\leq\Delta_{t}(\bm{v}_{t,i}|\langle S_{t,i-1}\rangle)+\delta$”; by the property (i) of Proposition 1, we have that, for any $\langle S_{t,i-1}\rangle\sim\langle\bm{S}_{t,i-1}\rangle$, $A(\langle S_{t,i-1}\rangle)$ is true with probability at least $1-\xi$, and then, its complementary event $\overline{A}(\langle S_{t,i-1}\rangle)$ is false with probability at most $\xi$. Thus, we have
$\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\max_{v\in V}\Delta_{t}(v|\langle\bm{S}_{t,i-1}\rangle)\right]$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\mathbb{P}[A(\langle\bm{S}_{t,i-1}\rangle)]\cdot\mathbb{E}_{\bm{v}_{t,i}}\left[\max_{v\in V}\Delta_{t}(v|\langle\bm{S}_{t,i-1}\rangle)|A(\langle\bm{S}_{t,i-1}\rangle)\right]\right]$
$\quad\quad+\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\mathbb{P}[\overline{A}(\langle\bm{S}_{t,i-1}\rangle)]\cdot\mathbb{E}_{\bm{v}_{t,i}}\left[\max_{v\in V}\Delta_{t}(v|\langle\bm{S}_{t,i-1}\rangle)|\overline{A}(\langle\bm{S}_{t,i-1}\rangle)\right]\right]$
$\leq\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\mathbb{P}[A(\langle\bm{S}_{t,i-1}\rangle)]\cdot\mathbb{E}_{\bm{v}_{t,i}}\left[\Delta_{t}(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta|A(\langle\bm{S}_{t,i-1}\rangle)\right]\right]$
$\quad\quad+\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\mathbb{P}[\overline{A}(\langle\bm{S}_{t,i-1}\rangle)]\cdot\mathbb{E}_{\bm{v}_{t,i}}\left[\Lambda|\overline{A}(\langle\bm{S}_{t,i-1}\rangle)\right]\right]$ (53)
$\leq\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\mathbb{E}_{\bm{v}_{t,i}}\left[\Delta_{t}(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta|\langle\bm{S}_{t,i-1}\rangle\right]\right]+\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\xi\cdot\Lambda\right]$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\mathbb{E}_{\bm{v}_{t,i}}\left[\Delta_{t}(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\right]|\langle\bm{S}_{t,i-1}\rangle\right]+\delta+\xi\Lambda$
$=\Delta_{t,i}+\delta+\xi\Lambda,$ (54)
where (53) holds since $\Lambda$, by definition, is an upper bound on the maximum increment achievable from the selection of any item. Furthermore, for any $h\in[b]$, let $\bm{v}^{*}_{t,h}$ be the $h$-th (random) item selected by the optimal adaptive policy $\pi^{*}_{t}$, let $\bm{S}^{*}_{t,h-1}:=\{\bm{v}^{*}_{t,1},\ldots,\bm{v}^{*}_{t,h-1}\}$ denote the set of the first $h-1$ items selected by $\pi^{*}_{t}$ and let $\Delta^{*}_{t,h}$ be the expected increment of $f_{t}$ due to the addition of item $\bm{v}^{*}_{t,h}$ to set $\bm{S}_{t,i-1}\cup\bm{S}^{*}_{t,h-1}$. Given $h\in[b]$, we have
$\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\max_{v\in V}\Delta_{t}(v|\langle\bm{S}_{t,{i-1}}\rangle)\right]$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\cup\bm{S}^{*}_{t,h-1}\rangle}\left[\max_{v\in V}\Delta_{t}(v|\langle\bm{S}_{t,{i-1}}\rangle)\right]$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\cup\bm{S}^{*}_{t,h-1}\rangle}\left[\mathbb{E}_{\bm{v}^{*}_{t,h}}\left[\max_{v\in V}\Delta_{t}(v|\langle\bm{S}_{t,{i-1}}\rangle)\right]\right]$
$\geq\mathbb{E}_{\langle\bm{S}_{t,i-1}\cup\bm{S}^{*}_{t,h-1}\rangle}\left[\mathbb{E}_{\bm{v}^{*}_{t,h}}\left[\Delta_{t}(\bm{v}^{*}_{t,h}|\langle\bm{S}_{t,{i-1}}\rangle)\right]\right]$
$\geq\mathbb{E}_{\langle\bm{S}_{t,i-1}\cup\bm{S}_{t,h-1}^{*}\rangle}\left[\mathbb{E}_{\bm{v}_{t,h}^{*}}\left[\Delta_{t}(\bm{v}_{t,h}^{*}|\langle\bm{S}_{t,i-1}\cup\bm{S}_{t,h-1}^{*}\rangle)\right]\right]$ (55)
$=\Delta^{*}_{t,h},$ (56)
where (55) holds by the adaptive submodularity (as $\langle\bm{S}_{t,i-1}\rangle\prec\langle\bm{S}_{t,i-1}\cup\bm{S}^{*}_{t,h-1}\rangle$). By combining (54) and (56) we obtain
$\Delta^{*}_{t,h}\leq\mathbb{E}_{\bm{S}_{t,i-1}}\left[\max_{v\in V}\Delta_{t}(v|\bm{S}_{t,{i-1}})\right]\leq\Delta_{t,i}+\delta+\xi\Lambda$ (57)
for any $h\in[b]$. Thus, by (57) we obtain
$GR^{*}_{t}(i-1,b)-GR_{t}(i-1)=\sum_{h\in[b]}\Delta^{*}_{t,h}\leq b(\Delta_{t,i}+\delta+\xi\Lambda)=b(GR_{t}(i)-GR_{t}(i-1)+\delta+\xi\Lambda),$
where the first equality holds since $GR^{*}_{t}(i-1,b)-GR_{t}(i-1)$ is the overall expected increment achieved by adding the $b$ items selected by $\pi^{*}_{t}$ to those already selected by ${\sf SingleGr}_{\delta,\xi}$. Finally, by rearranging the above inequality, we get inequality (52).
By combining (51) and (52) we obtain
$GR_{t}(i)\geq\frac{GR^{*}_{t}(i-1,b)}{b}+\left(1-\frac{1}{b}\right)GR_{t}(i-1)-(\delta+\xi\Lambda)\geq\frac{OPT_{t}(b)}{b}+\left(1-\frac{1}{b}\right)GR_{t}(i-1)-(\delta+\xi\Lambda).$ (58)
By applying (58) recursively from $i=1$ to $i=b$, with the base step $GR_{t}(0)=0$, we obtain
$GR_{t}(b)$ $\geq OPT_{t}(b)\left(\frac{1}{b}\right)\sum_{i=0}^{b-1}\left(1-\frac{1}{b}\right)^{b}-\sum_{i=0}^{b-1}\left(1-\frac{1}{b}\right)^{b}(\delta+\xi\Lambda)$
$\geq OPT_{t}(b)\left(\frac{1}{b}\right)\sum_{i=0}^{b-1}\left(1-\frac{1}{b}\right)^{b}-b(\delta+\xi\Lambda)$
$=OPT_{t}(b)\left(1-\left(1-\frac{1}{b}\right)^{b}\right)-b(\delta+\xi\Lambda)$
$\geq OPT_{t}(b)\left(1-\frac{1}{e}\right)-b(\delta+\xi\Lambda),$
where the last inequality follows from the known inequality $(1+x)^{r}<e^{rx}$, holding for any $x\geq-1$ and $r>0$, applied with $x=-1/B$ and $r=B$. By the above inequalities, the claim follows. ∎
Lemma 6.
For any vector $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b_{T}^{\prime})$ of non-negative integers such that $\sum_{t\in[T]}b_{t}^{\prime}=B$ we have
$GR(\vec{b}^{\prime})-(3\delta+2\xi\Lambda)B\leq GR(\vec{b}),$
where $\vec{b}$ is the vector returned by ${\sf BudgetGr}_{\delta,\epsilon}$.
Proof.
For any $t\in[T],i\in[n]$, let $\Delta_{t,i}$ denote the expected increment of the greedy single-round policy ${\sf SingleGr}_{\delta,\xi}$ with budget $n$ at round $t$, achieved when selecting the $i$-th item. By exploiting the adaptive submodularity as in (57), we obtain
$\Delta_{t,j}+\delta+\xi\Lambda\geq\Delta_{t,i}$ (59)
for any $i,j\in[n]$ with $i\leq j$ (that is, the approximate version of inequality (22) of Lemma 4).
Let $\tilde{\bm{\Delta}}_{t,i}$ be the estimation of $\Delta_{t,i}$ considered in ${\sf BudgetGr}_{\delta,\epsilon}$ and computed by Oracle2, so that the event $A:=$$|\tilde{\bm{\Delta}}_{t,i}-{\Delta}_{t,i}|\leq\delta$ for any $t\in[T]$ and $i\in[n]$” is true with probability at least $1-\xi$, and let $\overline{\bm{\Delta}}_{t,i}:=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$. Given a realization $(\tilde{\Delta}_{t,i})_{t,i}\sim(\tilde{\bm{\Delta}}_{t,i})_{t,i}$ such that event $A$ is true, $t\in[T]$, $i\in[n]$ and $i^{*}\in\arg\min_{i^{\prime}\in[i]}\tilde{\Delta}_{t,i^{\prime}}$ (that is, $\tilde{\Delta}_{t,i^{*}}=\overline{\Delta}_{t,i}$), we have
${\Delta}_{t,i}\leq\Delta_{t,i^{*}}+\delta+\xi\Lambda\leq\tilde{\Delta}_{t,i^{*}}+\delta+\delta+\xi\Lambda=\overline{\Delta}_{t,i}+2\delta+\xi\Lambda,$ (60)
where the first inequality holds by (59) and the second one holds since event $A$ is true by assumption; furthermore, we have
$\overline{\Delta}_{t,i}\leq\tilde{\Delta}_{t,i}\leq\Delta_{t,i}+\delta,$ (61)
where the second inequality again holds since event $A$ is true by assumption.
Now, let $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b_{T}^{\prime})$ be an arbitrary vector of non-negative integers such that $\sum_{t\in[T]}b_{t}^{\prime}$. As $\overline{\Delta}_{t,i}$ is non-decreasing in $i$, by using the same proof arguments as in Lemma 4, we have that the vector $\vec{b}$ returned by ${\sf BudgetGr}_{\delta,\epsilon}$ satisfies
$\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\Delta}_{t,i}\leq\sum_{t\in[T]}\sum_{i\in[b_{t}]}\overline{\Delta}_{t,i}.$ (62)
By exploiting (60), (61), (62) and the definition of event $A$, we have [^7] We observe that $(\Delta_{t,i})_{t,i}$ is a constant that does not depend on the realization of random variable $(\tilde{\bm{\Delta}}_{t,i})_{t,i}$.
$GR(\vec{b}^{\prime})$ $=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}$
$=\mathbb{P}[A]\cdot\mathbb{E}_{(\overline{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}\Bigg{|}A\right]+(1-\mathbb{P}[A])\cdot\mathbb{E}_{(\tilde{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}\Bigg{|}\overline{A}\right]$
$\leq\mathbb{P}[A]\cdot\mathbb{E}_{(\overline{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}(\overline{\bm{\Delta}}_{t,i}+2\delta+\xi\Lambda)\Bigg{|}A\right]+\xi\cdot\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Lambda$ (63)
$=\mathbb{P}[A]\cdot\mathbb{E}_{(\overline{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}(\overline{\bm{\Delta}}_{t,i})\Bigg{|}A\right]+B(2\delta+\xi\Lambda)+\xi B\Lambda$
$\leq\mathbb{P}[A]\cdot\mathbb{E}_{(\overline{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}]}(\overline{\bm{\Delta}}_{t,i})\Bigg{|}A\right]+B(2\delta+\xi\Lambda)+\xi B\Lambda$ (64)
$\leq\mathbb{P}[A]\cdot\mathbb{E}_{(\overline{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}]}(\Delta_{t,i}+\delta)\Bigg{|}A\right]+B(2\delta+\xi\Lambda)+\xi B\Lambda$ (65)
$=\mathbb{P}[A]\cdot\mathbb{E}_{(\overline{\bm{\Delta}}_{t,i})_{t,i}}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}]}\Delta_{t,i}\Bigg{|}A\right]+B\delta+B(2\delta+\xi\Lambda)+\xi B\Lambda$
$\leq\sum_{t\in[T]}\sum_{i\in[b_{t}]}\Delta_{t,i}+B\delta+B(2\delta+\xi\Lambda)+\xi B\Lambda$
$=GR(\vec{b})+B(3\delta+2\xi\Lambda)$ (66)
where $\overline{A}$ denotes the complementary event of $A$, (63) holds by (60) and $1-\mathbb{P}[A]\leq\xi$, (64) follows from (62), and (65) holds by (61). By (66), the claim follows. ∎
Lemma 7.
We have $\lambda\leq OPT$.
Proof.
By definition of $\lambda$, there exist $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$. As $OPT$ denotes the optimal adaptive value, we necessarily have $\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]\leq OPT$. Thus, by combining the obtained inequalities, the claim of the lemma follows. ∎